write an equation of a line with a slope -4/5 and a y-intercept of 4
Answer by mananth(16946)

m= - 4/ 5

Plug value of the slope and point ( 0,4 ) in
Y = m x + b
4.00 = -0 + b
b= 4 - -0
b= 4
So the equation will be
Y = - 4/5 x + 4
m.ananth@hotmail.ca